Leah and Ryanne went for a bike ride into the country. They averaged 24 miles per hour on the way out and 18 miles per hour on the way back, by the same roads. Their total travel time was 7 hours. How far did they go?
Answer by checkley71(8403)

X/24+X/18=7
(4X+3X)/72=7
7X/72=7 NOW CROSS MULTIPLY.
7X=7*72
X=7*72//7
X=72 MILES FOR THE TOTAL TRIP.
PROOF
72/24+72/18=7
3+4=7
7=7